Solve the problem with complete solutions. What will be the final temperature if a 4 x 10^7 J of heat is transferred to a 1 kg block of aluminum initially at 20 degree Celsius?


1
Expert's answer
2021-12-30T11:54:36-0500

The final temperature will be


"Q=cm(t_f-t_i),\\\space\\
t_f=t_i+\dfrac Q{cm}=20+\dfrac{4·10^7}{890·1}=44964°С."

This is too much, therefore, the aluminum will be the boiling point of 2470°C.